# Autoformalizing and Simulating Game-Theoretic Scenarios using LLM-augmented Agents

### v1

@@ 3.2.4. Constraint checking
valid_pd_payoffs(T,R,P,S,C,D):-
payoff(C,C,R,R),
payoff(C,D,S,T),
T>R,
payoff(D,C,T,S),
payoff(D,D,P,P),
R>P,
@@
Since the auto-formalization module does not impose constraints on the names of actions, this predicate also identifies which actions correspond to specific roles (e.g., determining that ‘coop’ = C).
3.3. Autoformalization Module
Algorithm 1 Generating game rules predicates from natural language descriptions Mensfelt et al. (2024). PD stands for Prisoner’s Dilemma.
@@ 3.3. Autoformalization Module
To perform autoformalization, a Large Language Model is employed. The framework supports the use of any language model by providing an abstract interface. Autoformalization of a natural language description of a game or strategy is achieved through one-shot prompting, using game-dependent predicates for the Prisoner’s Dilemma and the tit-for-tat strategy as examples. After generating the code, its syntactic correctness is validated. If errors are detected, the LLM receives feedback, including the erroneous lines, error messages, and instructions for correcting the predicates. The max_attempts parameter specifies the maximum number of correction attempts. The overview of autoformalization algorithm is presented in Listing 1.
3.4. Simulation
In our previous work Mensfelt et al. (2024), we manually validated the semantic correctness of generated programs. In this study, we introduce a tournament-based mechanism to automate the semantic correctness validation for scenarios and strategies where a target total payoff or a payoff matrix is available. In this approach, once a scenario is autoformalized, a copy of an agent is created and supplied with a specified strategy. The agent and its copy then play a specified number of rounds (clones mode). Comparing the agent’s total payoff to the target total payoff provides an approximate validation. When a ground truth payoff matrix is available, it enables exact semantic validation by verifying whether the payoff received for each unique combination of moves matches the target payoff for that combination. Additionally, we perform constraint checking to ensure that the payoff matrix satisfies the specified constraints for a specified game.
The tournament can also evaluate which strategy is most successful, on average, within a given game scenario across a specified number of rounds. This is achieved by supplying agent copies with different strategies and having them compete in a round-robin mode, where each agent plays against every other.
The main parameters of a tournament are as follows:
  • game: can be provided by predefined rules or natural language description to autoformalize,
  • strategies: can be provided as a list of predefined strategy rules or a list of natural language descriptions to autoformalize,
  • mode: playing against a single clone or round-robin,
  • agents number,
  • rounds number.
There are two methods for determining the winners of the tournament:
  1. (1)
    Selecting agents which total payoff matches their target payoff – this is applicable if a target payoff is provided for each agent.
  2. (2)
    Selecting agents that achieved the highest total payoff.
Both methods were utilized in the evaluation of the framework, as detailed in the following sections.
4. Methods
4.1. Parameters
Common parameters
LLM GPT-4o
@@ 4.1. Parameters
maximum output tokens $1024$
@@
Table 3. Common experimental parameters.
To evaluate the framework, we conducted three experiments: Experiment 1 tested the autoformalization of game scenarios, Experiment 2 employed the autoformalized game rules in tournament’s round-robin mode, and Experiment 3 examined the autoformalization of strategies. Table 3 provides the experimental parameters shared across all experiments, while Table 4 details the parameters unique to each specific experiment.
Parameter Exp. 1 Exp. 2 Exp. 3
games $55$ $5$ $1$
agents per game $5$ $1$ $5$
strategies 1 $6$ $5$
rounds $4$ $10$ $4$
mode clones round-robin clones
Table 4. Experiment-specific parameters.
4.2. Exp. 1: Autoformalization of Game Descriptions
In the first experiment, we used a dataset of 55 natural language game-theoretic scenarios to evaluate the syntactic and semantic accuracy of logic programs generated by agents. The dataset included 5 scenarios that employed common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games introduced in Section 2.1. The remaining 50 scenarios – 10 for each game – used alternative descriptions that avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma). For each scenario, five agents were generated with autoformalized game rules and then engaged in tournament, using the tit-for-tat strategy against a clone employing the anti-tit-for-tat strategy, testing all four possible pairs of moves in four rounds. Each agent had a maximum of five attempts to produce syntactically correct code; agents failing after five attempts were labelled as syntactically incorrect. Post-tournament, each agent’s individual and total payoffs were compared to the target, with matches indicating semantic correctness.
4.3. Exp. 2: Axelrod’s Tournament
@@
Table 5. Strategies utilised in the evaluation.
In the second experiment, we selected five agents created in Experiment 1, one for each of the five games. For each game, we conducted a tournament based on the design of Axelrod’s tournament Axelrod and Hamilton (1981). In this setup, six copies of each agent were created and assigned one of six strategies; each agent then played against every other agent, including itself. The strategies used are listed in Table 5.
To adapt the strategies to games without explicit cooperation and defection concepts, we defined a “default move” and its “opposite” as the basis for certain strategies. The default moves were: “Cooperate” for the Prisoner’s Dilemma, “Dove” for Hawk-Dove, “Stag” for Stag Hunt, “Opera” for the Battle of the Sexes, and “Heads” for Matching Pennies.
The main metric for evaluating strategy performance was the total payoff achieved by an agent using a given strategy in each game. To allow for cross-game comparisons of strategy effectiveness, total payoffs were normalized with the following formula:
@@
We then evaluated the semantic correctness of the autoformalized strategies. To achieve this, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as a reference example. For the other five strategies, the autoformalization module was provided with their natural language descriptions, similar to those shown in Table 5.
@@
Each autoformalized strategy was then assigned to an agent with predefined Prisoner’s Dilemma game rules. This agent played four rounds against a clone utilizing the anti-tit-for-tat strategy. The agent’s total payoff was compared to a target payoff specific to each strategy to assess correctness. For the random strategy, correctness was also manually validated. This tournament setup was repeated five times for each strategy to ensure consistency in results.
5. Results and Discussion
5.1. Exp. 1: Autoformalization of Game Descriptions
Attempts 1 2 3 4 5
Count 93% 7% 0 0 12%
Table 6. Distribution of autoformalization attempts number for 275 agents.
Table 6 shows the distribution of attempts needed to generate syntactically correct code for the 275 agents created in the first experiment. For 93% of agents, syntactically correct code was generated on the first attempt. When the correct code was not generated initially, 7% of agents succeeded on the second attempt, while 12% could not produce correct code within the allowed attempts. This latter case highlights a need for further refinement of the feedback loop to improve success rates. However, the examination of the non-aggregated distribution of the invalid agents showed that an agent without syntactically valid game program was generated in no more than 1 in 5 instances.
Figure 2. The percentage of syntactic and semantic correctness of generated game rules for each game.
The overall syntactic correctness rate was 96%, while semantic correctness reached 87%, closely aligning with the results of our previous work (98% and 88%, respectively). Figure 2 depicts the distribution of syntactic and semantic correctness across various games. The Battle of the Sexes and Stag Hunt emerged as the most error-prone scenarios, exhibiting the lowest semantic correctness rates, though still exceeding 80%.
The total-payoff-based approximate validation demonstrated 88% semantic correctness, while the exact individual-payoff-based validation achieved 87%. Five instances were incorrectly labelled as semantically correct by the approximate method, indicating that the results of the approximate and exact methods were closely aligned. A comparison of constraint checking and exact semantic validation revealed a few cases where the autoformalized payoff matrix satisfied the constraints, but errors in other parts of the program – such as role assignment – resulted in incorrect payoff assignments. This demonstrates that exact validation is feasible only when the ground truth payoff matrix is available.
5.2. Exp. 2: Axelrod’s Tournament
Figure 3. Normalized total payoff for each strategy by game.
@@ 5. Results and Discussion
Figure 3 presents the normalized total payoffs for each strategy across all games, while Table 7 provides the average normalized total payoff for each strategy across games. Overall, the best-response strategy proved to be the most effective on average, followed closely by tit-for-tat. Best-response was most effective in all games except the Battle of the Sexes and the Prisoner’s Dilemma. In the Prisoner’s Dilemma, best-response achieved a lower payoff than the anti-default-move strategy (analogous to “defect”), as it initiated with the “cooperate” move. The results demonstrate that agents with autoformalized rules can effectively simulate and assess the performance of various strategies within different game-theoretic scenarios.
Strategy Payoff [%]
anti-default-move 50.90
anti-tit-for-tat 55.22
best-response 84.08
default-move 73.96
random 61.08
tit-for-tat 79.84
Table 7. Average normalized total payoffs for each strategy.
5.3. Exp. 3: Autoformalization of Strategies
Strategy Correctness [%]
anti-default-move 100.00
anti-tit-for-tat 100.00
best-response 40.00
default-move 100.00
random 60.00
Table 8. Percentage of correctness for each strategy.
Table 8 presents the percentage of correctly autoformalized strategies. The simpler strategies—default-move, anti-default-move, and anti-tit-for-tat—were always correctly autoformalized. In contrast, the random strategy achieved correct autoformalization in 60% of cases, and best-response, which requires selecting the move that maximizes payoff in response to the opponent’s previous move, was correctly autoformalized in 40% of cases. Since all strategies were generated using only the tit-for-tat implementation as an example, increasing the diversity of examples provided to the autoformalization module could potentially improve its accuracy in correctly formalising more complex strategies.
6. Related Work
@@
Autoformalization Wu et al. (2022), has shown promising results in translating natural language into mathematical formalisms, enabling the integration of large language models with formal tools for solving reasoning tasksWu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022). This approach has also been successfully extended to creating logical representations from natural language descriptions. For instance, Yang et al. Yang et al. (2023) applied GPT-3 with few-shot prompting to translate natural language sentences into formal representations, which were then inputted into answer set programs. While effective on NLP benchmarks, this method faced occasional translation errors. To further enhance performance, Pan et al.Pan et al. (2023) combined in-context learning with self-refinement based on solver evaluations, surpassing standard LLM and chain-of-thought prompting approaches on logical reasoning benchmarks. In the context of temporal logic, Cosler et al.Cosler et al. (2023) addressed the problem of translation errors with an interactive system allowing users to refine sub-translations. Fine-tuning LLMs has also proven effective for boosting accuracy in translations to temporal logicsChen et al. (2023) and deductive datasets Feng et al. (2023).
7. Conclusions and Future Work
@@
We have presented a framework for the autoformalization of game-theoretic scenarios, specifically those modelled by $2\times 2$ simultaneous-move games. The framework also enables autoformalization of strategies for these games and the simulation of tournaments based on the autoformalized rules. The generated programs undergo three levels of validation: syntactic validation by consulting a solver, functional testing through gameplay, and semantic validation – either approximate or exact – by comparing achieved payoffs with target outcomes. Evaluation of the framework demonstrated high syntactic and semantic accuracy of the code generated by agents across various interaction scenarios.
@@
Future work will focus on extending the framework to model games beyond $2\times 2$ simultaneous-move scenarios. Our previous work Mensfelt et al. (2024) demonstrated that our sample solver for the Prisoner’s Dilemma could successfully serve to generate code for structurally similar games, such as rock-paper-scissors (with a different number of moves) and sequential versions of the Prisoner’s Dilemma. However, achieving successful generalization to a broader class of games will require assembling a more diverse set of in-context learning examples and developing methods to retrieve the most relevant examples for autoformalization in specific scenarios. The flexibility and generalisability of the introduced framework also make it applicable to games and game-like interactions beyond game theory.
Another direction for future work involves enhancing the feedback loop to more effectively leverage a large language model. At present, the LLM receives feedback solely from the trace generated when consulting the solver with the produced code. In the next phase, we plan to expand this feedback mechanism by incorporating runtime errors, thereby creating a more robust autoformalization module. Finally, in cases where unsupervised autoformalization proves infeasible for complex scenarios, we plan to introduce an interactive mode. This mode will allow users to iteratively refine the generated code alongside the autoformalization module. These interactively generated programs will then be added to our existing program library, expanding the resources available for in-context learning and enabling the autoformalization of increasingly complex scenarios.
{acks}

### v2

@@ 3.2.2 Validity checking
valid_pd_payoffs(T,R,P,S,C,D) if
payoff(C,C,R,R) and
payoff(C,D,S,T) and
T>R and
payoff(D,C,T,S) and
payoff(D,D,P,P) and
R>P and
@@
Since the autoformalization module does not impose constraints on the names of actions, this predicate also identifies which actions names correspond to specific actions.
3.3 Autoformalization Module
Algorithm 1 Generating game rules predicates from natural language descriptions. PD stands for Prisoner’s Dilemma.
@@ 3.3 Autoformalization Module
To perform autoformalization, a large language model is employed. The framework supports the use of any language model by providing an abstract interface. Autoformalization of a natural language description of a game or strategy is achieved through one-shot prompting, using game-dependent predicates for the Prisoner’s Dilemma or the tit-for-tat strategy as examples, respectively. After generating the code, its syntactic correctness is validated. If errors are detected, the LLM receives feedback, including the erroneous lines, error messages, and instructions for correcting the predicates. The max_attempts parameter specifies the maximum number of correction attempts. The overview of autoformalization algorithm is presented in Listing 1 Mensfelt et al. (2024).
3.4 Simulation
The framework is employed in this work in two primary modes: validation and simulation. These modes facilitate the evaluation of both the semantic correctness of generated programs and the effectiveness of strategies within various game scenarios.
In validation mode, the framework ensures the semantic correctness of generated programs by comparing their performance against predefined criteria. Once a scenario is autoformalized, an agent is created and provided with a specified strategy. The agent then plays against a clone of itself for a set number of rounds. If a ground truth payoff matrix is available, the framework performs semantic validation by verifying that the payoff for each unique combination of moves matches the target payoff for that combination. Additionally, if the ground truth type of the game is known, the framework enforces constraint checking to confirm that the payoff matrix adheres to all specified constraints for the given game.
In simulation mode, the framework evaluates and compares strategies to determine their effectiveness within a game scenario. Agents equipped with different strategies compete in a round-robin tournament, where each agent plays against every other. By simulating multiple rounds, the framework identifies the strategy that performs best on average, offering insights into the relative strengths and weaknesses of competing approaches.
The parameters of the tournament are as follows:
  • Agent Pool: Contains all agents participating in the tournament. Each agent can be equipped with game rules and strategies either through autoformalization or predefined configurations.
  • Number of Rounds: Specifies how many rounds each agent will play.
  • Match Maker: A function that determines the pairing of agents for matches.
  • Target Payoffs: Optional parameter used for semantic validation.
The framework provides two methods for determining the tournament’s outcome:
  1. 1.
    Selecting agents whose total payoff matches their target payoff, applicable when target payoffs are specified (validation).
  2. 2.
    Selecting agents that achieve the highest total payoff, suitable for assessing the success of strategies (simulation).
Furthermore, the customizable Match Maker function allows for experiments with tailored matching protocols
4 Methods
4.1 Parameters
Parameter Value
LLM Claude 3.5 Sonnet
GPT-4o
@@ 4.1 Parameters
maximum output tokens $2048$
@@
Table 2: Common experimental parameters.
Parameter Exp. 1 Exp. 2 Exp. 3 Exp. 4
games $55$ $55$ $5$ $5$
agents in game $5$ $5$ $1$ $1$
strategies $1$ $1$ $6$ $6$
rounds $4$ $4$ $10$ $10$
match maker S S RR S
Table 3: Experiment-specific parameters. S stands for a mode in which an agent plays against its clone with an opposite strategy, while RR stands for round robin.
To evaluate the framework, we conducted four experiments: i) autoformalization of game scenarios with numerical (Exp. 1) and ii) non-numerical (Exp. 2) payoffs, iii) using the autoformalized game rules in a round-robin tournament (Exp.3), and iv) autoformalization of strategies (Exp. 4). Table 2 provides the experimental parameters shared across all experiments, while Table 3 details the parameters unique to each specific experiment.
4.2 Autoformalization of Interaction Descriptions (Numerical Payoffs)
We used a dataset of 55 natural language game-theoretic scenarios with numerical payoffs to evaluate the syntactic and semantic accuracy of generated logic programs. The dataset included 5 scenarios that employed common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games introduced in Section 2.1. The remaining 50 scenarios—10 for each game—used alternative descriptions that avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma). For each scenario, five agents autoformalized game rules and then played tournament against their clones, generating all four possible pairs of moves in four rounds. Each agent had a maximum of five attempts to produce syntactically correct code; agents failing after five attempts were labelled as syntactically incorrect. Post-tournament, each agent’s individual and total payoffs were compared to the target, with matches indicating semantic correctness.
4.3 Autoformalization of Interaction Descriptions (Non-numerical Payoffs)
This experiment was analogous to the previous one, but it utilized interaction scenarios without numerical payoffs (e.g., ”If both remain silent, they get minor sentences,” to describe the outcome of mutual cooperation in PD). Therefore, the autoformalizing LLM had to infer the appropriate payoffs correctly. Since it was not possible to provide a ground truth matrix in this case, we validated the semantic correctness by performing constraint checks (described in Sect. 3.2) to determine whether the payoff matrix satisfied the constraints for the ground truth game type (Sect. 2.1).
4.4 Axelrod’s Tournament
@@
Table 4: Strategies utilized in the evaluation.
To demonstrate the framework’s usefulness in analysing interaction scenarios, we selected five agents created during the first experiment, one for each of the five games. For each game, we conducted a tournament modelled after Axelrod’s tournament Axelrod and Hamilton (1981). In this setup, six copies of each agent were created and assigned one of six strategies. Each agent then played against every other agent, including itself. The strategies used are listed in Table 4.
The main metric for evaluating strategy performance was the total payoff achieved by an agent using a given strategy in each game. To allow for cross-game comparisons of strategy effectiveness, total payoffs were normalized using:
@@ 4.5 Autoformalization of Strategies
We then evaluated the semantic correctness of the autoformalized strategies. To achieve this, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as an example. For the other five strategies, the autoformalization module was provided with their natural language descriptions, similar to those shown in Table 4.
@@
Each autoformalized strategy was then assigned to an agent with predefined Prisoner’s Dilemma game rules. This agent played four rounds against a clone utilizing the anti-tit-for-tat strategy. The agent’s total payoff was compared to a target payoff specific to each strategy to assess correctness. For the random strategy, correctness was also manually validated. This tournament setup was repeated five times.
5 Results and Discussion
5.1 Autoformalization of Interaction Descriptions
1 2 3 4 5
Numeric payoffs
Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
GPT-4o 96.77% 2.87% 0.00% 0.00% 0.36%
Non-numeric payoffs
Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
GPT-4o 97.64% 2.02% 0.34% 0.00% 0.00%
Table 5: Distribution of attempts at creating syntactically correct code for each model.
Table 5 shows the distribution of attempts needed to generate syntactically correct code for both models. The type of description (numeric vs. non-numeric) had no influence on syntactic correctness of the generated code. Claude 3.5 Sonnet achieved syntactically correct code in 100% of cases and GPT-4o in nearly 100%. The few errors observed involved non-Prolog comment characters and non-escaped special characters in strings, all of which were successfully corrected on the second attempt. This demonstrates that providing a specific line that generated the error in the feedback prompt significantly enhances an LLM’s ability to self-correct compared to merely providing an error message Mensfelt et al. (2024). The only case where GPT-4o was unable to correct a syntactic error involved the error message “Unexpected end of file“ which indicated the end of the file as the error’s source. However, the actual line generating the error was not provided in the feedback prompt.
Syntactic Runtime Semantic
Numeric payoffs
Claude 3.5 Sonnet 100.0% 88.0% 86.0%
GPT-4o 100.0% 90.0% 86.0%
Non-numeric payoffs
Claude 3.5 Sonnet 100.0% 79.0% 67.0%
GPT-4o 100.0% 91.0% 68.0%
Table 6: Syntactic, runtime, and semantic correctness for both models and types of scenario descriptions.
The type of description had no influence on runtime errors in GPT-4o (runtime correctness of 90% and 91% for numeric and non-numeric payoffs, respectively, Table 6). However, it did have an influence in the case of Claude 3.5 Sonnet, with a 9 percentage point decrease to 79% for non-numerical payoffs. A runtime error in this context means that the program was syntactically correct; however, while querying a predicate during runtime—most often to calculate payoffs—a predicate necessary to execute the query was missing. This highlights the need to extend the autoformalization feedback loop to the runtime stage and leverage the tracing mechanism of the solver to identify and create missing predicates.
Numeric Payoffs
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 89.1% 96.4% 68.5% 96.4% 87.3%
Semantic 89.1% 96.4% 68.5% 89.1% 87.3%
Non-Numeric Payoffs
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 82.1% 85.7% 74.0% 70.7% 83.1%
Semantic 82.1% 77.8% 34.2% 64.0% 83.1%
Table 7: Correctness by game for Claude 3.5 Sonnet for game descriptions with numeric and non-numeric payoffs.
Numeric Payoffs
BS HD MP PD SH
Syntactic 98.3% 100.0% 100.0% 100.0% 100.0%
Runtime 83.1% 89.1% 96.4% 89.1% 90.9%
Semantic 79.7% 89.1% 90.9% 85.5% 83.6%
Non-Numeric Payoffs
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 93.2% 88.1% 87.3% 96.5% 91.5%
Semantic 93.2% 67.8% 30.2% 73.7% 76.3%
Table 8: Correctness by game for GPT-4o for game descriptions with numeric and non-numeric payoffs.
The semantic correctness was lower than runtime correctness for both models and description types (Table 6). This means that even though an agent was able to select a move during runtime and update its state, the semantics of the program were not correct. For descriptions with numerical payoffs, this implies that the payoff matrix does not contain the required payoffs. For descriptions without numerical payoffs, it indicates that the payoff matrix does not satisfy the constraints of the required game. Tables 7 and 8 present the syntactic, runtime, and semantic correctness for different games.
The runtime and semantic correctness varied depending on the model and the game. Notably, Matching Pennies consistently resulted in low semantic correctness for non-numerical descriptions, with 34.2% correctness for Claude 3.5 Sonnet and 30.2% correctness for GPT-4o. This may be attributed to some descriptions specifying which player wins if the choices match or mismatch, without explicitly stating that the other player will lose—a constraint of the game. Extending the autoformalization feedback loop to include semantic evaluation has the potential to significantly improve correctness.
5.2 Axelrod’s Tournament
Figure 2: Normalized total payoff for each strategy by game.
@@ 5.2 Axelrod’s Tournament
Figure 2 presents the normalized total payoffs for each strategy across all games. Overall, the best-response strategy proved to be the most effective on average, followed closely by tit-for-tat. Best-response was most effective in all games except the Battle of the Sexes and the Prisoner’s Dilemma. In the Prisoner’s Dilemma, best-response achieved a lower payoff than the anti-default-move strategy (analogous to “defect”), as it initiated with the “cooperate” move. The results demonstrate that agents with autoformalized rules can effectively simulate and assess the performance of various strategies within different modelled scenarios.
5.3 Autoformalization of Strategies
Strategy Claude 3.5 Sonnet GPT-4o
anti-default-move 100.00% 100.00%
anti-tit-for-tat 100.00% 100.00%
best-response 80.00% 100.00%
default-move 100.00% 100.00%
random 80.00% 100.00%
Table 9: Percentage of semantic correctness for each strategy.
Table 9 presents the percentage of correctly autoformalized strategies. GPT-4o achieved 100% correctness in autoformalizing all five strategies. Claude 3.5 Sonnet produced one incorrect strategy for best-response and one for random, which are the strategies that deviate the most from the tit-for-tat strategy presented as an example in the prompt. One error occurred during runtime, and another during semantic validation. Overall, the high correctness demonstrates that the GAMA framework has the potential to effectively autoformalize both game rules and gameplay strategies.
6 Related Work
@@
Autoformalization Wu et al. (2022), has shown promising results in translating natural language into mathematical formalisms, enabling the integration of large language models with formal tools for solving reasoning tasks Wu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022). This approach has also been successfully extended to creating logical representations from natural language descriptions. For instance, Yang et al. (2023) applied GPT-3 with few-shot prompting to translate natural language sentences into formal representations, which were then inputted into answer set programs. While effective on NLP benchmarks, this method faced occasional translation errors. To further enhance performance, Pan et al. (2023) combined in-context learning with self-refinement based on solver evaluations, surpassing standard LLM and chain-of-thought prompting approaches on logical reasoning benchmarks. In the context of temporal logic, Cosler et al. (2023) addressed the problem of translation errors with an interactive system allowing users to refine sub-translations. Fine-tuning LLMs has also proven effective for boosting accuracy in translations to temporal logics Chen et al. (2023) and deductive datasets Feng et al. (2023).
7 Conclusions and Future Work
@@
We have presented a framework for the autoformalization of interaction scenarios, specifically those modelled by $2\times 2$ simultaneous-move games. The framework enables the autoformalization of strategies for these games and the simulation of tournaments based on the autoformalized rules. The generated programs undergo three levels of validation: syntactic validation by consulting a solver, functional testing through gameplay, and semantic validation by comparing achieved payoffs with target outcomes or by checking the game constraints. Evaluation of the framework demonstrated high syntactic and semantic accuracy in the code generated by agents across various interaction scenarios.
@@
Future work will focus on extending the framework to model games beyond $2\times 2$ simultaneous-move scenarios. Our previous work Mensfelt et al. (2024) demonstrated that using our sample solver for the Prisoner’s Dilemma we could successfully generate code for structurally similar games, such as rock-paper-scissors (with a different number of moves) and sequential versions of the PD. However, achieving successful generalization to a broader class of games will require assembling a more diverse set of in-context learning examples and developing methods to retrieve the most relevant examples for autoformalization in specific scenarios. Another direction for future work involves extending the autoformalization feedback loop to include runtime and semantic validation checks, further improving the correctness of the generated programs.
These extensions will provide a robust framework for generating reasoning modules for decision-making agents in a wider class of games and game-like interactions.
Acknowledgments